A Visit to the 23rd International İstanbul Film Festival

David Pierce

2004.04.27; minor corrections, 2004.08.07

On Wednesday, April 21, 2004, Ayşe and I caught a 13.30 bus from Ankara’s autogare (in Turkish phonetic spelling, otogar). Rain struck the windshield of the bus for most of the six-hour trip to İstanbul, but the wind was not strong, and the Bosphorus was calm when we took the ferry to the European side.

Rather than stay with friends living some fifteen kilometers from the festival cinemas, we got a hotel room near the action. We had dinner at our favorite vegetarian restaurant, Zencefil (whose name means ‘ginger’). This has moved across the street to a new and larger site with a garden.

It was good to have slept locally before heading out Thursday morning for a day of watching movies. Last year we took an overnight train, but I don’t sleep well on those, one reason perhaps being that the beds are not quite so long as I am.

We saw four movies on Thursday, three on Friday, and four on Saturday. We could have seen four on Friday, but instead we had dinner with the friends we weren’t staying with. Technically we could have seen five movies on any given day: the screenings were at 11.00, 13.30, 16.00, 19.00 and 21.30. But the last one was too late for us. The movies were shown on five screens in four different movie-houses, but these are all within a block of each other on the pedestrianized İstiklal Caddesi (Independence Avenue); nearby is a cafeteria where even vegetarian cinephiles can grab a good quick meal. (One cinema on the Asian side of İstanbul also shows festival movies.)

Here are the movies we saw, in order. All movies are from 2003 or 2004; we didn’t aim to see any oldies this time, though I noticed that Citizen Kane was sold out.

1. Work Hard, Play Hard, Jean-Marc Martout (France and Belgium). The director’s first, and one of the best we saw. The theme is what capitalism can do to your human sympathies. Young Philippe works for a management consulting firm in the towers of La Défense outside Paris. On the Métro, he scolds a man who is feeling up an attractive young woman. Philippe ends up dating this woman, Eva. Meanwhile, his job is to improve the efficiency of an aluminium firm. The workers aren’t told that the point is to make the firm more attractive to a buyer; but they figure it out, and they figure that a number of them will be fired. Philippe almost quits his own job, saying he can’t be responsible for the workers’ losing theirs. Eva has suggested as much. But Philippe’s boss persuades him to continue his task.

The English title of the movie is actually the motto of Philippe’s company. Their job is to induce other people to work hard and productively. Their reward is to play hard at a company party—which Philippe attends rather than stay home with Eva and her sick child. The child wasn’t really sick; Eva was just testing Philippe to see what was important to him.

Nobody in the movie spoke the equivalent of Wall Street’s ‘You want a friend, get a dog!’ Nobody was shown getting filthy rich through exploitation of the workers. Philippe observes that a worker has to spend a lot of time walking between two machines that could have been closer together. ‘Didn’t you ever suggest moving the machines?’ asks Philippe. ‘No’ says the worker. But the machines probably should be closer together; Philippe is right.

Now, it seems to me that the best way to bring about such efficiencies is to have the workers sharing ownership in the factory; but the movie itself doesn’t make such radical suggestions. The movie seems a fair portrait of what goes on—though perhaps I’m not one to know, really, since I have never worked for a corporation.

2. The Professional, Dusan Kovacevic, Serbia-Montenegro. It ended up winning a jury prize in the festival competition. I think the prize was deserved, though I was bothered by the sometimes-silly treatment of a serious business; also I didn’t always follow the politics. Workers at a Belgrade printing plant are preparing to go on strike while the director is in his large office making out with his secretary. But the director is not so bad; as an opponent of Milosevic, he was trailed unwittingly by a spy for ten years. He learns this, because the spy visits him to confess. So we get flashbacks as the spy recounts the man’s history.

Silliness arises when, for example, the two man are raising their shirts and lowering their pants to compare scars from a botched assassination attempt, by car-crash, on the one by the other. Naturally the secretary picks that moment to walk in, and she faints.

3. Fuse, Pjer Zalica, Bosnia-Herzegovina. The catalogue also lists Austria, Turkey and France as having contributed to the film. A town in ‘Muslim’ Bosnia is near the border with the Serbian section. Word comes that President Clinton will be visiting; so the people had better display lots of brotherhood with the Serbians. A Nato commander is there to ensure this. A sort of Potemkin village is set up under his nose: Old records in the mayor’s office are burned; Serbs who fled the town are paid to return to their old homes for a day. But the forcible reintegration of peoples is not entirely unsuccessful.

One amusing moment of many: the townspeople sew a big American flag, but in ignorance they make the stars a Communist red.

More seriously, one old man talks to the ghost of his dead son. He is sure his son is alive somewhere, as some Americans believe that PoWs are still alive in Vietnam. But like Reagan’s would-be assassin, he also gets the idea that a violent act will bring attention to his cause…

I enjoyed the last movie more than the previous; but the people seemed pretty juvenile in both. Finally for Day 1, we saw:

4. Small Town, Jan Kraus, Czech Republic. Supposedly, most Czech critics didn’t like it, because of its unpleasant picture of life in the transition to capitalism—or life in general in the places referred to by the title. I can’t describe the movie well: there were too many people, with interrelations not entirely clear. In the beginning, before he passes out drunk, a Communist party official has requested a wife, or at least a ‘crotch’ as the subtitle would have it. So the townspeople find a local woman willing to service the fat slob. Later, after free elections, it seems that the ‘Erotic Party’ has won something; and there is a bizarre scene of cheerleaders running across tables and over beer-mugs as seated men look under skirts.

People are supposed to be running businesses now, but they don’t really know how. Police investigate blood, only to find that a man is slaughtering pigs in his basement, since his other business isn’t working out.

I can only recall vignettes like this; I can’t thread them together, if that is even possible. The film was not boring; I don’t suppose it could be boring when there was a strip-tease show at the the high-school, dutifully attended by townspeople of all (reasonably mature) ages and both sexes: the naked performer led men onstage so they could eat the banana stuck between her legs as their wives and children watched. But I didn’t exactly see the point.

Our day 2 at the festival was really day 14 of the festival’s 16. We saw:

5. Young Gods, Jukka-Pekka Siili, Finland. I wondered how this movie would avoid charges of child pornography. I suppose it does so by being reasonably discreet for its theme. Four male friends just out of high school form a club. Their purpose is to capture on videotape their sexacts with girls. One of the boys hasn’t had a girlfriend yet, so he needs some coaching. It turns out that there is a girl with her eye on him, but their assignation ends badly to say the least. Almost everything ends badly, one might say, except that it is good that secret taping of intimate activities is shown not to be all fun and games.

And there seem to be deeper things going on. The club’s leader, Taavi, is an orphan whose parents left him a fortune and a large collection of home videos. The boy claims he can’t really see things in the family mansion directly, but only in the little screen of the video camera. And the family video collection is missing a crucial volume, showing how the parents came to die; the boy obtains the tape from a reluctant trustee…

At the beginning of the movie, Taavi receives a surprise birthday present from his friends male and female: a public nude hug in downtown Helsinki. The youngsters are taken to the police station, but one boy among them talks the policewoman out of fining them. She’s young herself, and finds something attractive in the way the boy talks to her…

Frequently throughout the movie, the color images are replaced with black and white pictures from security cameras. We are all being watched.

The Finnish director of Young Gods spoke a few words at the screening. He said people in Finland either loved or hated the movie. We didn’t wait for him to say much else, as we figured he could do his best speaking through the movie itself, and anyway it was time to get some lunch before heading to another cinema and another movie:

6. Alila, Amos Gitai, Israel and France. The setting is an apartment complex in Tel-Aviv. A contractor named Ezra is building some kind of addition with Chinese laborers. Ezra’s son is awol from his military service. Ezra seems still to love his ex-wife, who however has another man living with her (a man I first thought was another son, actually). Yet another man rents an apartment in the complex so he can have a place to make love to his mistress; but her screams of pleasure annoy the neighbors. Another woman in the complex uses ‘Arab’ as a term of abuse, as in ‘We’re not Arabs, why don’t you have coffee with us?’ She is extremely unattractive as a person; but nobody in the movie is all that likable. Ezra’s awol son might be the most sympathetic, if only he could enunciate his beliefs. Ezra disowns him for his lack of patriotism, but later relents.

A holocaust survivor in the complex has a Chinese maid. She speaks English, though she seems to understand some things in Hebrew. Yet, as she is washing dishes, a news broadcast (in Hebrew) on the radio notes that a bomb has gone off in a district where many foreigners congregate; her expression doesn’t change as she switches to a music station. The scene is creepy, and is another reason why the movie is not that pleasant to watch. But like all the movies I’m talking about, this one shows why a film industry in every country is valuable, and why the tendency of American movies to put these local industries out of business is a bad thing.

The next movie was a gem:

7. Since Otar Left, Julie Bertucelli, France and Belgium. Those countries provided the director and the funding, but the setting is mostly Georgia. The movie is another director’s first. Three generations of women live in an apartment in Tbilisi. Otar is the son/brother/uncle, but he lives in Paris. The family have been Francophiles since before 1917. They rely on money sent back by Otar, though they have a dacha in the country and a large library of leather-bound French volumes that they could sell if they were desperate. Otar’s mother, Eka, says dear old Stalin would straighten out the economy if he were around, and by the way, he never ordered anybody’s death. Eka dotes on her absent son, to the consternation of her daughter, Marina—whom Eka scolds for not treating her daughter, Ada, correctly.

Otar had been a medical student in Moscow; now he is working illegally in construction in Paris. We never see him before word comes to Tbilisi that he has died in an accident at a worksite. The news is kept from his mother; his niece writes letters to her grandmother in her uncle’s name, saying that his working hours no longer permit him to telephone.

I can’t generally approve of keeping secrets. Indeed, Ada grows uncomfortable: she observes that her mother is simply afraid that Otar will become an absolute saint in Eka’s eyes if she knows that he is dead. In short, Otar’s sister is acting selfishly.

The truth does come out, sort of: I’ll only say that it comes out in a satisfactory way.

There’s a very nice scene when Marina and Ada are away at the dacha, and Eka goes by herself to an amusement park, buys two cigarettes, and smokes them on a ride that takes her above the green trees.

Melda and Mustafa came with us to see Otar before we all went to eat dinner. Melda’s ancestors are from Georgia, and she says her family keep secrets as in the movie. But I’ve observed such secret-keeping elsewhere in Turkey, and it must happen all over the world. A big Turkish industrialist died recently of a cancer that he didn’t know he was dying of, because his family kept the news from him. (One wonders if they didn’t want him to rearrange his finances, though I think Turkey doesn’t give one much control over one’s estate after one dies.)

Day 3:

8. Vodka Lemon, Hiner Saleem, Armenia et al. Another gem. The setting is rural Armenia in winter. The director, originally from Iraq, is Kurdish, as are most of the people in the movie. Mainly the movie is a slowly blossoming love-affair between a widower, Hamo, and a widow, Nina, who come to be aware of each other during their daily visits to the local snow-covered cemetery. To the driver of the rickety bus that takes them there, Nina keeps saying that she will pay tomorrow. Hamo periodically sells his old belongings for a few dollars; some of these he finally uses to pay Nina’s bus fare. Nina has a job, but only at a stall selling bottles of ‘vodka lemon’ to passing truck-drivers. ‘Why does it taste like almonds?’ asks a customer. ‘This is Armenia’ she says. On a good day, 17 bottles are sold, but it’s not enough for the owner to keep the stand open. Hamo wishes the Russians were still in power; the people may not have had freedom then, but they had everything else.

The people seem to be neither Christian nor Muslim. There is a sheep sacrifice for a colorful wedding that is held out in the snow. The bride is Hamo’s grand-daughter; the groom reneges on the bride-price, which included a job for the bride’s father in Novosibirsk.

People don’t seem to mind much whether they are indoors or out; they take their household chairs out and sit around in the snow drinking vodka. The snow is beautiful and covers up hillsides that, like the ones I saw in Eastern Turkey last September, must be bereft almost even of grass.

9. Parallel Journeys, Derviş Zaim and Panicos Chrysanthou, Turkey and Greece. The catalogue lists the directors as Turkish and Greek, but I think they are both Cypriots. We saw this documentary on the day when Cyprus voted on the UN peace plan. Supposed Communists were demonstrating on İstiklal Caddesi against the plan. They may say that they don’t like the EU, but I think they can fairly be called Turkish imperialists. In fact the Turkish Cypriots voted mostly for the peace plan, and the Greek Cypriots voted mostly against.

As for the documentary, it showed people recalling the Turkish invasion and talking about the other side. Most people said they could live with people from the other community, although one young teacher praised Atatürk and said she didn’t want the communities to become mixed. One Turkish man recalled how he escaped the extermination of his village, presumably by Greeks, whereas one Greek man recalled being treated reasonably fairly by Turkish soldiers. Yet this brings out an annoying feature of the documentary: do the stories of these two men leave one with an accurate summary of what really happened in 1974? At the beginning, the filmmakers give a very brief historical summary; then they are silent, and I don’t think they should be.

10. Last Life in the Universe, Pen-ek Ratanaruang, Thailand. This may be the best Thai film in years, they say; never having seen a Thai film, I wouldn’t know. Neither would I expect the film to appeal to the masses of filmgoers in Thailand or anywhere else—and indeed the festival organizers put the film in the ‘Mined Zone’. We had bad experiences in the Mined Zone last year, but this movie was quite lovely. There are gunfights at the beginning and end; but the movie’s Australian cinematographer Christopher Doyle was on hand afterwards to explain, to a complaining woman in the audience, that perhaps such action scenes were required by the movie’s backers.

The leading man is Japanese and works in the library at the Japan Cultural Centre in Bangkok. He is what some would call anal-retentive: when he arranges cans of beer in his refrigerator, he turns them so the labels all face the same way. He also plays at committing suicide, because death would be so relaxing; but something always interrupts.

He ends up having to flee Bangkok, and he holes up in the large isolated house of a Thai woman whose dirty dishes are stacked on the sofa and everywhere else. They try to communicate in each other’s native tongue, but mostly they have to resort to English. As in Vodka Lemon, slowly something develops. He washes her dishes; she thanks him…

The cinematography is supposed to be great. I’m not sure exactly that this means, but the movie is very pleasant to watch. There is something fascinating in the empty pool surrounded by overgrown grass behind the woman’s house.

11. The Agronomist, Jonathan Demme, USA. I didn’t see the director’s Silence of the Lambs (or anything else). The man named in the title of this documentary is Haitian patriot Jean Dominique, whose Radio Haiti Inter earned the hatred of the ruling class and eventually Dominique’s assassination. Meanwhile, Demme had known Dominique for years and had filmed him in several places, in Haiti or in exile in New York. He was a wonderful speaker, intense and animated, but always cheerful— chuckling as he recalled going to prison. Dominique was light-skinned and evidently had wealth, but as far as I could tell he worked only for the common people. He would broadcast news of the fall of dictators elsewhere, so that the people might think ‘Hey, Papa Doc or Baby Doc or Raoul Cedras could go like them!’ He supported Aristide, but then asked why Aristide was making deals with rich people. He observed that the US could stop the suffering in Haiti with a phone call.

And that was it for our movies. We caught a bus home on Sunday morning. During the journey, I started reading Why People Hate America (which I thought could have been better done).